f(x)=x^2-2, then f(x+2)-f(x)=?????
Answer by stanbon(48569)

f(x)=x^2-2, then f(x+2)-f(x)
----------
f(x+2) = (x+2)^2-2
= x^2+4x+2
--------
f(x)=x^2-2
----------
Subtract f(x) from f(x+2) to get:
f(x+2)-f(x) = 4x+4
==============
Cheers,
Stan H.